OpenStudy (anonymous):

How do u solve this: Find the roots of the quadratic equation x2 − 8x + 8 = 0 . x1 = x2 =

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (anonymous):

a=1 , b= -8 , c=8 usding the equation x1= 4+2sqrt2 x2= 4-2sqrt2

OpenStudy (anonymous):

so we just plug it into the quadratic formula and it gives us those two answers?

OpenStudy (anonymous):

yes. \[ax^2 + bx + c = 0 \implies x = \frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

thanks guys

OpenStudy (radar):

Car2021 your equation can be factored giving you; (x-8)(x-1)

OpenStudy (radar):

Since this product is 0 you know that the factors are zero (at least one or both) x=8 and x=1

OpenStudy (anonymous):

@radar i don't think it can be factrorized, try to spread what you type u will get x^2 -9x+8

OpenStudy (radar):

Hey funky girl you are so right and that does not equal the equation. Please ignore Car2021 and thanks for straightening me out lol

OpenStudy (anonymous):

You are welcome =)